Frank Natterer

Schriftenverzeichnis

1. Einschließungen für die großen Eigenwerte bei gewöhnlichen Differentialgleichungen zweiter und vierter Ordnung. Numer . Math. 13, 78-93 (1969).

2. Numerische Behandlung singulärer Sturm-Liouville-Probleme. Numer. Math. 13, 434-447 (1969).

3. Numerical methods of high-order accuracy for singular nonlinear boundary value problems. (Zusammen mit P.G. Ciarlet und R.S. Varga). Numer. Math. 15, 87-99 (1970).

4. Schranken für die Eigenwerte gewöhnlicher Differentialgleichungen durch Spline- Approximation. Numer. Math. 14, 346-354 (1970).

5. Berechnung der Norm der Inversen eines Differentialoperators zweiter Ordnung. Computing 5, 214-220 (1970).

6. Schranken für die Norm der Greenschen Funktion gewöhnlicher Differentialoperatoren. Numer. Math. 17, 402-408 (1971).

7. A generalized spline method for singular boundary value problems of ordinary differential equations. Lin. Alg. Appl. 7, 189-216 (1973).

8. Eine Erweiterung des Maximumprinzips für den Laplaceschen Operator. (Zusammen mit B. Werner). Numer. Math. 22, 149-156 (1974).

9. Das Randwertproblem für eine Klasse singulärer gewöhnlicher Differentialausdrücke. Aequationes Mathematicae 12, 207-228 (1975).

10. Das Differenzenverfahren für singuläre Rand-Eigenwertaufgaben gewöhnlicher Differentialgleichungen. Numer. Math. 23, 387-409 (1975).

11. Berechenbare Fehlerschranken für die Methode der Finiten Elemente. ISNM 28, 109-128 (1975).

12. Über die punktweise Konvergenz Finiter Elemente. Numer. Math. 25, 67-77 (1975).

13. Optimale L_{2}-Konvergenz Finiter Elemente bei Variationsungleichungen. Bonner Mathematische Schriften 89, 1-11 (1976).

14. Uniform convergence of Galerkin′s method for splines on highly nonuniform meshes. Math. Comp. 31, 457-468 (1977).

15. The finite element method for ill-posed problems. RAIRO Analyse numerique 11, 271-278 (1977).

16. Regularisierung schlecht gestellter Probleme durch Projektionsverfahren. Numer. Math. 28, 329-341 (1977).

17. Berechnung oberer Schranken für die Norm der Ritz-Projektion auf Finite Elemente. ISNM 39, 236-245 (1978).

18. Acceleration of convergence for finite element solutions of the Poisson equation on irregular meshes. (Zusammen mit A. Louis). TR Fachbereich 10 der Universität des Saarlandes.

19. Numerical inversion of the Radon transform. Numer. Math. 30, 81-91 (1978).

20. Numerical analysis of ill-posed problems. Fundacao Universidade de Brasilia, Trabalho de Matemática No. 145 (1979).

21. A Sobolev space analysis of picture reconstruction. SIAM J. Appl. Math. 39, 402-411 (1980).

22. Efficient implementation of “optimal” algorithms in computerized tomography. Math. Meth. Appl. Sci. 2, 545-555 (1980).

23. On the smoothing property of the Bayes approach to picture reconstruction. TR Fachbereich 10 der Universität des Saarlandes (1978).

24. An ultra-fast algorithm in tomography. (Zusammen mit K.H. Löw). TR Fachbereich 10 der Universität des Saarlandes (1981).

25. On the inversion of the attenuated Radon transform. Numer. Math. 32, 431-438 (1979).

26. Vorausberechnung des Grubenklimas in sonderbewetterten Betrieben. (Zusammen mit Quack, H. und Güsmann, B.). Abschlußbericht über das Projekt Nr. 01 VB 285 - Z 13 TAP 0005 des BMFT, Januar 1979.

27. Genauigkeitsfragen bei der numerischen Rekonstruktion von Bildern, in Ansorge, R. (ed.): Vorträge auf der Tagung über Numerische Mathematik aus Anlaß der Emeritierung von L. Collatz, ISNM Birkhäuser Verlag, Januar 1979.

28. The ill-posedness of Radon′s integral equation. Proc. of the Symposion on ill-posed problems: Theory and Practice, Delaware, USA, October 2-6 (1979).

29. The identification problem in emission computed tomography. In: Herman, G.T. - Natterer, F. (eds.): Mathematical Aspects of Computerized Tomography. Springer 1981.

30. On the order of regularization methods. ISNM 63, 189-203 (1983).

31. Computerized tomography with unknown sources. SIAM J. Appl. Math. 43, 1201-1212 (1983).

32. Error bounds for Tikhonov regularization in Hilbert scales. Appl. Anal. 18, 29-37 (1984).

33. Exploiting the ranges of Radon transforms in computerized tomography. In: Deuflhard, P. - Hairer, E. (eds.): Numerical Treatment of Inverse Problems in Differential and Integral Equations. Birkhäuser 1983.

34. Some non-standard Radon problems. In: Boffi, V. - Neunzert, H. (eds.): Applications of Mathematics in Technology. Teubner 1984.

35. Mathematical problems of computerized tomography. (Invited paper, zusammen mit A. Louis). Proceedings of the IEEE 71, 379-389 (1983).

36. Einige Beiträge der Mathematik zur Computer-Tomographie. ZAMM 64, T252-T260 (1984).

37. Discretization ill-posed problems. Publicazioni dell′istituto di analisi globale e applicazioni, serie Problemi non ben posti ed inversi, n. 9, Firenze 1983.

38. Some ill-posed problems arising in connection with Radon′s integral equation. Publicazioni dell′istituto di analisi globale e applicazioni, serie Problemi non ben posti ed inversi, n. 8, Firenze 1983.

39. Efficient evaluation of oversampled functions. J. Computational and Applied Mathematics 14, 303-309 (1986).

40. Computable error bounds for linear systems via backward analysis. Preprint, Institut für Numerische und instrumentelle Mathematik, Münster 1985.

41. Fourier reconstruction in tomography. Numer. Math., 47, 343-353 (1985).

42. Numerik des Radonschen Problems. Jber. d. Dt. Math. Verein. 87, 108-126 (1985).

43. Numerical treatment of ill-posed problems. In: Talenti, G. (ed.): Inverse problems, p. 142-167. Springer 1986.

44. The Mathematics of Computerized Tomography. Wiley - Teubner 1986, 222 S.

45. Severely ill-posed Radon problems. In: Engl, H.W. - Groetsch, C.W. (eds.): Inverse and ill-posed problems, p. 247-259. Academic Press 1987.

46. Attenuation correction in emission tomography. In: Sabatier, P. (ed.): Inverse Problems, p. 21-33. Academic Press 1987.

47. Mathematics and tomography. In: McKenna, J. - Temam, R. (eds.): Proceedings of the first international conference on industrial and applied mathematics, SIAM, Philadelphia 1988, p. 183-199.

48. Regularization techniques in medical imaging. In: Viergever, M.A. - Todd-Pokropek, A. (eds.): Mathematics and Computer Science in Medical Imaging, p. 127-142. NATO ASI series 1987.

49. The attenuated Radon transform. In: Viergever, M.A. - Todd-Pokropek, A. (eds.): Mathematics and Computer Science in Medical Imaging, p. 271-278. NATO ASI series 1987.

50. Resolution in computerized tomography. In: Meyer-Ebrecht, D. (Hrsg.): ASST′87, p. 29-36. Springer 1987.

51. Identifying a coefficient in a stationary Boltzman equation. Publicationi dell′istituto di analisi globale e applicazioni, Problemi non ben posti ed inversi, n. 38. Firenze 1988.

52. Attenuation correction in PET. (Zusammen mit H. Herzog, KfA Jülich), Mathematical Methods in the Applied Sciences 15, p. 321-330 (1992).

53. An inverse problem for the transport equation and integral geometry. Proceedings of the AMS conference on Integral Geometry, Arcata 1989, Contemporay Mathematics, AMS 1990.

54. Tomography with unknown directions. To appear in the Proceedings of the Oberwolfach Conference on “Mathematical Methods in Tomography”, 1990. Springer 1991.

55. A discrete Gelfand-Levitan theory, Preprint.

56. A Lanczos type method for inverse Sturm-Liouville problems. Proceedings of the Centre for Mathematics and its Applications, Australien National University, 31, 82-88, (1992).

57. Discrete inverse evolution problems. In: Sabatier, P. (ed.): Inverse Problems in Action, p. 29-37. Springer 1990.

58. 2D sampling in tomography. In: Laurent, P.J. - Le Méhauté, A. - Schumaker, L.L. (eds.): Curves and Surfaces, Academic Press 1991.

59. Basic algorithms in tomography. (Zusammen mit A. Faridani), in: Grünbaum, F.A. - Helton, J.W. - Khargonekar, P. (eds.): Signal Processing PART II: Control Theory and Applications, p. 321-334, Springer 1990.

60. A tutorial in tomography. Procedings of the conference on “Problèmes Inverses”, Rocquencourt 1990. INRIA, Domaine de Voluceau BP 105, F-78153 Le Chesnay Cedex.

61. Local and global tomography. (Zusammen mit A. Faridani, F. Keinert, E.L. Ritman, K.T. Smith), in: Grünbaum, F.A. Helton, J.W. - Khargonekar, P. (eds.): Signal Processing PART II: Control Theory and Applications, p. 241-256, Springer 1990.

62. Mathematische Methoden der Computer–Tomographie. Rheinisch- Westfälische Akademie der Wissenschaften, Vorträge N375, S. 7-25. Westdeutscher Verlag 1990.

63. The Mathematics of Computerized Tomography (Russisch). Erweiterte Ausgabe von 44. MIR, Moskau 1990.

64. A finite difference method for the inverse scattering problem at fixed frequency (with F. Wübbeling). Proceedings of the Lapland conference on Inverse Problems, June 1992.

65. Sampling in fan beam tomography, SIAM J. Appl. Math., 53, 358-380 (1993).

66. Determination of tissue attenuation in emission tomography of optically dense media, Inverse Problems 9, p. 731-736 (1993).

67. Efficient scanning in fan beam tomography. SPIE 1767, 320-325 (1992).

68. Tomography with few data, in: Anger et al. (eds.): Inverse Problems: Principle and Applications in Geophysics, Technology, and Medicine. Akademie-Verlag 1993.

69. Sampling and resolution in CT, in: M.M. Lavrent’ev (ed.): Computerized Tomography, p. 343-354. VSP Utrecht, 1995.

70. Recent developments in x-ray tomography, in: Quinto, E.T. - Cheney, M. - Kuchment, P. (eds.): Tomography, Impedance Imaging, and Integral Geometry. Lecture Notes in Applied Mathematics, Vol. 30, AMS 1994.

71. Reconstruction of Process Tomography Images (Keynote Review), in: Beck, M.S. et al. (eds.): Proceedings of ECAPT’94, March 24-26, Porto 1994.

72. Finite Difference Methods for Inverse Problems, Proceedings of the Workshop on Inverse Problems, HoChiMinh City, January 17-19 (1995).

73. A Propagation-Backpropagation Method for Ultrasound Tomography, Inverse Problems 11, 1225-1232 (1995).

74. Numerical Solution of Bilinear Inverse Problems, Technical Report 19/96-N, Fachbereich Mathematik der Universität Münster.

75. Accurate Attenuation Correction in Spect Without Transmission Measurements, (with A. Welch, R. Clack, G.T. Gullberg). IEEE NSS-MIC Conference, San Franzisco, Oct. 21-28 (1995).

76. An initial value approach for inverse Helmholtz problems, Proceedings of the Conference “Inverse Problems in Medical Imaging and Nondestructive Testing”, Oberwolfach, February 4-10, 1996, 1996. To appear Springer 1996.

77. Algorithms in Tomography. Proceedings of the Conference ‘State of the Art in Numerical Analysis”, York, April 1-4, 1996. To appear Oxford University Press 1996.

78. An Algorithm for 3D Ultrasound Tomography, Talk to be presented at the Conference “Inverse problems of wave propagation and diffraction”, Aix-les-Bains, September 23-27, 1996.

79. Resolution and Reconstruction for a Helical CT-Scanner, Technical Report 20/96-N, Fachbereich Mathematik der Universität Münster.

0LaTeX: berg f berger listenfn natlit1.tex